# chessformer.bib

### v1

@@ Abstract
Transformer models have demonstrated impressive capabilities when trained at scale, excelling at difficult cognitive tasks requiring complex reasoning and rational decision-making. In this paper, we explore the application of transformer models to chess, focusing on the critical role of the position encoding within the attention mechanism. We show that in chess, transformers endowed with a sufficiently versatile position encoding can match existing chess-playing models at a fraction of the computational cost. Our architecture significantly outperforms AlphaZero at 8x fewer FLOPS and matches prior grandmaster-level transformer-based agents at 30x fewer FLOPS.
1 Introduction
@@
Chess has long been regarded as a proving ground for advancing artificial intelligence (AI). The importance of strategic planning and rational decision-making in strong chess play make the game an ideal domain in which to test systems aiming to master these human-like cognitive capabilities.
@@
Traditional chess engines employ a specialized tree-search algorithm paired with a handcrafted evaluation function. Most modern chess engines follow this recipe but use an efficiently updated neural network (NNUE) rather than a handcrafted evaluation function.
@@
The AlphaZero engine [AlphaZero] introduced a different recipe based on Monte Carlo Tree Search (MCTS) and deep neural networks. The open-source Leela Chess Zero (Lc0) recreated this approach, building on the same principles with several subsequent enhancements. These engines use large networks that predict not only a position evaluation but also a policy distribution over subsequent moves which is used to guide the search process.
AlphaZero used a convolution-based residual network, which was the leading architecture of the time. However, the state of the art in AI has since shifted from convolution-based residual networks to attention-based transformer models. Transformers resolve the issue of small receptive fields that limit convolution-based architectures and demonstrate "LLM-like scaling", exhibiting impressive capabilities when trained at scale.
@@
We demonstrate that the effectiveness of transformers in chess depends heavily on the choice of position encoding in the attention mechanism. Based on this observation, we adopt the versatile position encoding scheme of Shaw et al. [shaw2018selfattention] and train models endowed with this technique and other enhancements at scale, calling the resulting architecture the ChessFormer. This architecture substantially outperforms prior work in playing strength and puzzle-solving ability at a fraction of the computational cost.
Our contributions are as follows:
@@
  • We present a simple yet performant “ChessFormer” architecture for transformers in chess and apply this architecture across model scales.
  • We ablate the position encoding of our architecture to motivate design choices.
@@
  • We provide a detailed comparison of our models against prior work to demonstrate the superiority of our techniques.
@@
  • We open-source our training code at https://github.com/Ergodice/lczero-training.
@@
This paper is organized as follows. Section 2 describes the vanilla self-attention formulation and compare three position encodings of varying versatility. Section 3 describes our training setup. We compare the results of our models to prior work and run ablation experiments in Section 4. Related work is reviewed in Section 5, and concluding remarks are presented in Section 6.
@@
Figure 1: Elo strength comparison of agents constructed from our CF-6M and CF-240M models against prior art. Our evaluation methodology is described in Section 4.
2 Background
2.1 Self-Attention
@@
Given a sequence of tokens $(\mathbf{x}_{1},\ldots\mathbf{x}_{n})$ where $\mathbf{x}_{i}\in\mathbb{R}^{d}$, self-attention returns a sequence $(\mathbf{z}_{1},\ldots\mathbf{z}_{n})$ where $\mathbf{z}_{i}\in\mathbb{R}^{d}$. Using projection matrices $W^{Q}$, $W^{V}$, and $W^{K}$ in $\mathbb{R}^{d\times d}$, a logit $e_{ij}$ for each pair of tokens $(i,j)$ is computed using scaled dot-product attention:
@@
The attention weights are obtained via softmax:
@@
Finally, the output is computed as a weighted sum of value projections:
@@
2.2 Position Encodings
Because self-attention is permutation-invariant, positional information must be introduced to the model through some kind of position encoding. Many position encoding techniques, like rotary position embeddings [rope], are designed to decay the attention values with the Euclidean distance between tokens.
While Euclidean distance is a useful inductive bias for language and vision applications, we conjecture that it is not sufficient to capture the topology of the chessboard. A square with a bishop, for example, might need to attend to squares on the same diagonal rather than squares that are nearby. To test this, we compare several position encodings. Changes to the vanilla self-attention formulation described in Section 2.1 are marked in blue.
The simplest form of position encoding, the absolute position embedding, adds a possibly learnable bias to each token before the attention layer. Calling these biases $(\mathbf{c}_{1},\ldots\mathbf{c}_{n})$, the absolute position embedding modifies the formula for attention logits as follows:
$e_{ij}=\frac{((\mathbf{x}_{i}{\color[rgb]{0,0,1}+\mathbf{c}_{i}})W^{Q})((\mathbf{x}_{j}{\color[rgb]{0,0,1}+\mathbf{c}_{j}})W^{K})^{T}}{\sqrt{d}}$ (4)
Unlike absolute position embeddings, relative position encodings model the positions of tokens relative to each other. One simple variant introduces relative biases $d_{ij}$ which are added to the attention logits:
@@ 2.2 Position Encodings
For one-dimensional inputs, these bias terms are shared among pairs of tokens $(i,j)$ and $(k,l)$ for which $i-j=k-l$. We adopt the two-dimensional analog of this technique, where two pairs of tokens share a relative position encoding if their horizontal and vertical displacements are the same.
@@
A more general encoding introduced by Shaw et al. [shaw2018selfattention], which we adopt in our final architecture, models the positional relationship between tokens $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ by introducing learnable vectors $a_{ij}^{Q}$, $a_{ij}^{K}$, and $a_{ij}^{V}$ in $\mathbb{R}^{d}$. The calculation for attention logits is altered to:
$e_{ij}=\frac{(\mathbf{x}_{i}W^{Q}{\color[rgb]{0,0,1}+a_{ij}^{Q}})(\mathbf{x}_{j}W^{K}{\color[rgb]{0,0,1}+a_{ij}^{K}})^{T}}{\sqrt{d}}.$ (6)
@@
The output is updated to propagate $a_{ij}^{V}$ in addition to the output of the value projection:
$\mathbf{z}_{i}=\sum_{j=1}^{n}\alpha_{ij}(\mathbf{x}_{j}W^{V}{\color[rgb]{0,0,1}+a_{ij}^{V}}).$ (7)
@@
While this encoding is computationally expensive at large token counts, our models have a context length of 64, making the additional cost negligible.
3 Experimental Setup
We train all our models in the supervised setting on selfplay data generated via the AlphaZero process. The selfplay setup uses a transformer model of roughly 100 million parameters at 600 nodes per move. This model is almost identical in architecture to our main runs, although it uses a slightly weaker position encoding. All models and ablations of the ChessFormer architecture employ a standard encoder-only backbone with context length 64 with one of the position encodings described in Section 2. We describe the architecture in more detail in Appendix A and note other techniques which did not improve performance in Appendix D.
All models were trained with the Nadam optimizer with $\beta_{1}=0.9$, $\beta=0.98$, and $\epsilon=10^{-7}$, and gradient clipping was set to 10. The final model checkpoints we use for evaluation were generated with stochastic weight averaging [swa].
@@ 3 Experimental Setup
Our largest model, CF-240M, has 15 encoder layers with an embedding depth of 1024, 32 heads per layer, and a feedforward depth of 4096, for a total of 243 million parameters. It was trained for 3.7 million steps with a batch size of 4096 on 8 A100 GPUs using data parallelism. The learning rate was initialized at $1*10^{-3}$ and manually reduced to $3*10^{-4}$ at 3.2 million steps and $1*10^{-4}$ at 3.6 million steps. The dataset consisted of 500 million games generated from mid-2023 to mid-2024, with each game containing roughly 200 positions.
@@
We also train a smaller model, CF-6M, with 8 encoder layers, an embedding depth of 256, 8 heads per layer, and a feedforward depth of 256, for a total of 6 million parameters. The same configuration is used for ablation runs. CF-6M and its ablations were each trained on a single A100 GPU with a batch size of 2048. The learning rate was initialized at $5*10^{-4}$ and reduced to $1.58*10^{-4}$ and $5*10^{-5}$ at 1.6 and 1.8 million steps, respectively. The dataset consisted of 53 million games generated in April 2024.
Our training targets consist of a mix of auxiliary policy and value targets in addition to the main policy and value heads. More details can be found in Appendix B. None of our training runs exhibited overfitting, which is likely due to the size of the dataset relative to the parameter counts of our models.
4 Results
@@
We compare the playing and puzzle-solving abilities of agents constructed from our models to prior work and ablate the position encoding in our architecture to motivate design choices.
@@
We consider two types of agents: those that select the move which is ranked highest in the policy vector predicted by the model, and those that emulate a depth 1 search by evaluating the model for each legal move and selecting the move that maximizes the score. The policy strategy requires a single model evaluation while the score maximization strategy requires an evaluation for each legal move. To estimate the FLOPs used by an agent of the score maximization type, we multiply the model FLOPs by 20, which is roughly equal to the average number of legal moves available in a position.
@@
We construct both types of agents from our models. We denote an agent by its model name followed by “policy” or “value” to indicate whether the agent uses the predicted policy vector or the score maximization approach (e.g., CF-240M-policy). We construct two agents from the AlphaZero model in the same way. Additionally, our analysis includes models from Ruoss et al. [ruoss2024grandmasterlevel] which use the score maximization approach. We use the final checkpoints of their main runs having parameter counts of 9 million, 136 million, and 270 million, which we refer to as GC-9M, GC-136M, and GC-270M, respectively.
4.1 Playing Strength
@@
To estimate the playing strength of our agents, we play 1,000 games with them against the GC-270M agent of Ruoss et al. [ruoss2024grandmasterlevel], tying the Elo rating of GC-270M to the value reported in that paper. We use the same opening book as Ruoss et al. to achieve a similar comparison to theirs. Elo values for the GC and AlphaZero agents are taken from that paper. The results are shown in LABEL:table:main and Figure 1.
4.2 Puzzles
@@
In addition to playing strength, we also evaluate the puzzle-solving ability of our agents using the subset of the Lichess puzzle database curated by Ruoss et al. [ruoss2024grandmasterlevel]. Our evaluation follows the same methodology as theirs, with an agent deemed to solve a puzzle only if it correctly picks each move in the sequence of the solution. The results are shown in LABEL:table:main. In Figure 2, we also graph the accuracy of our two CF-240M agents on this test set against the GC-136M and GC-270M agents of Ruoss et al. [ruoss2024grandmasterlevel].
@@
Figure 2: Puzzle-solving ability on puzzles rated 1000-3000 of our CF-240M-policy and CF-240M-value agents against the GC-136M and GC-270M agents of Ruoss et al. [ruoss2024grandmasterlevel].
4.3 Ablations
@@
We train ablations of the CF-6M model, comparing the three position encodings described in Section 2. As shown in LABEL:table:ablate, our chosen encoding, that of Shaw et al. [shaw2018selfattention], substantially outperforms both the relative bias and absolute position encodings. The reported value loss and accuracy statistics are those of the main “result” value head. At this parameter count, we have observed that doubling model size yields an increase in policy accuracy of around 1.5%.
4.4 Analysis
@@
Our main results can be found in LABEL:table:main. Our agents consistently outperform prior work in both playing strength and puzzle-solving ability at a fraction of the computational cost. Our CF-6M-policy agent outperforms the AlphaZero-policy agent in both puzzle-solving ability and Elo performance at 8x fewer FLOPS. At 30x fewer FLOPS, our CF-240M-policy agent matches the puzzle accuracy and exceeds the Elo performance of the grandmaster-level agent GC-270M of Ruoss et al. [ruoss2024grandmasterlevel].
5 Related Work

### v2

# Mastering Chess with a Transformer Model

@@ Abstract
Transformer models have demonstrated impressive capabilities when trained at scale, excelling at difficult cognitive tasks requiring complex reasoning and rational decision-making. In this paper, we explore the application of transformers to chess, focusing on the critical role of the position representation within the attention mechanism. We show that transformers endowed with a sufficiently expressive position representation can match existing chess-playing models at a fraction of the computational cost. Our architecture, which we call the Chessformer, significantly outperforms AlphaZero in both playing strength and puzzle solving ability with 8x less computation and matches prior grandmaster-level transformer-based agents in those metrics with 30x less computation. Our models also display an understanding of chess dissimilar and orthogonal to that of top traditional engines, detecting high-level positional features like trapped pieces and fortresses that those engines struggle with. This work demonstrates that domain-specific enhancements can in large part replace the need for model scale, while also highlighting that deep learning can make strides even in areas dominated by search-based methods.
Index Terms:
Chess, AI, Transformer, Machine Learning
I Introduction
@@
Chess has long been regarded as a proving ground for artificial intelligence (AI). The importance of strategic planning and rational decision-making in strong chess play make the game an ideal domain in which to test systems aiming to master these human-like cognitive capabilities.
@@
Traditional chess engines employ a specialized tree-search algorithm paired with a handcrafted evaluation function. Most modern chess engines follow this scheme but use an efficiently updated neural network (NNUE) as their evaluation function.
@@
The AlphaZero engine [1] introduced a different recipe based on Monte Carlo Tree Search (MCTS) and deep neural networks. This approach was recreated by the open-source Leela Chess Zero (Lc0), which builds on the same principles with several subsequent enhancements. These engines use large networks that predict not only a position evaluation but also a policy vector, a distribution over subsequent moves which is used to guide the search process.
AlphaZero used a convolution-based residual network, which was the state of the art at the time. It was based on AlphaGo, which had prior achieved superhuman performance in Go [2]. However, as the authors of the AlphaZero paper note, convolution-based models may be poorly suited for chess. Long-range interactions feature much more prominently in chess than in Go, and convolutions are poorly equipped to deal with these because of their small receptive fields. Transformer models, on the other hand, are based on a global self-attention operation and resolve the issue of small receptive fields. They can also exhibit impressive abilities when trained at scale, powering large language models such as OpenAI’s GPT-4 [3].
@@
We show that the effectiveness of transformers in chess depends heavily on the choice of position representation in the attention mechanism. Among three representations of varying expressivity, we select for our final architecture the scheme of Shaw et al. [4], training models endowed with this technique and other enhancements at scale. We call the resulting architecture the Chessformer.
Our main contributions are:
@@
  • We present a simple yet performant architecture for transformers in chess and apply this architecture at scale.
  • We ablate the position representation in the attention mechanism of our models to demonstrate the criticality of a sufficiently expressive representation.
@@
  • We provide a detailed comparison of our models against prior work to demonstrate the superiority of our methods.
  • We analyze the attention maps and playstyle of our largest model, CF-240M.
@@
  • We open-source our training code. [^1] Available on Github at https://github.com/Ergodice/lczero-training.
@@
This paper is organized as follows. Section II goes over the vanilla self-attention formulation and compares three position representations of varying expressivity. Section III describes our training setup. Section IV compares our models to prior work and reports results for ablating the position representation. Section V analyzes the attention maps of our largest model, CF-240M, while Section VI describes its playstyle. Section VII reviews related work, and Section VIII provides concluding remarks.
@@
Figure 1: Elo strength by floating point operations per evaluation (FLOPS) of agents constructed from our CF-6M and CF-240M models (red) against prior art (blue). Our evaluation methodology is described in Section IV.
II Background
II-A Self-Attention
@@
Given a sequence of tokens $(\mathbf{x}_{1},\ldots\mathbf{x}_{n})$ where $\mathbf{x}_{i}\in\mathbb{R}^{d}$, self-attention returns a sequence $(\mathbf{z}_{1},\ldots\mathbf{z}_{n})$ where $\mathbf{z}_{i}\in\mathbb{R}^{d}$. Using projection matrices $W^{Q}$, $W^{V}$, and $W^{K}$ in $\mathbb{R}^{d\times d}$, a logit $e_{ij}$ for each pair of tokens $(i,j)$ is computed through scaled dot-product attention:
@@
Attention weights are obtained via softmax:
@@
Finally, the output is computed as a weighted sum of value information:
@@
II-B Position Representations
Because self-attention is permutation-invariant, positional information must be introduced to the model through some kind of position representation. Many position representation techniques, like rotary position embeddings [5], are designed to decay the attention weights with the Euclidean distance between tokens.
While Euclidean distance is a useful inductive bias for language and vision applications, it fails to capture the topology of the chessboard. A square with a bishop, for example, might need to attend to squares on the same diagonal rather than squares that are nearby. To demonstrate the need for more expressivity in the position representation, we compare three options: absolution position embeddings, relative biases, and the scheme of Shaw et al. [4]. Changes to the vanilla self-attention formulation described in Section II-A are marked in blue.
The simplest form of position representation, the absolute position embedding, adds a possibly learnable bias to each token prior to the attention layer. Calling these biases $(\mathbf{c}_{1},\ldots\mathbf{c}_{n})$, the absolute position embedding adds these biases to the $\mathbf{x}_{i}$ prior to the attention calculation:
$\mathbf{x}_{i}\mapsto\mathbf{x}_{i}+{\color[rgb]{0,0,1}\mathbf{c}_{i}}$ (4)
Unlike absolute position embeddings, relative position representations model the positions of tokens relative to each other. One simple variant introduces relative biases $d_{ij}$ which are added to the attention logits:
@@ II-B Position Representations
For one-dimensional inputs, these bias terms are shared among pairs of tokens $(i,j)$ and $(k,l)$ for which $i-j=k-l$. We adopt the two-dimensional analog of this technique, where two pairs of tokens share a relative bias if their horizontal and vertical displacements are the same.
@@
A more general position representation introduced by Shaw et al. [4], which we adopt in our final architecture, models the positional relationship between tokens $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ by introducing learnable vectors $a_{ij}^{Q}$, $a_{ij}^{K}$, and $a_{ij}^{V}$ in $\mathbb{R}^{d}$. The calculation for attention logits is altered to:
$e_{ij}=\frac{(\mathbf{x}_{i}W^{Q}{\color[rgb]{0,0,1}+a_{ij}^{Q}})(\mathbf{x}_{j}W^{K}{\color[rgb]{0,0,1}+a_{ij}^{K}})^{T}}{\sqrt{d}}$ (6)
@@
The output, meanwhile, is updated to propagate signal from $a_{ij}^{V}$ in addition to the output of the value projection:
$\mathbf{z}_{i}=\sum_{j=1}^{n}\alpha_{ij}(\mathbf{x}_{j}W^{V}{\color[rgb]{0,0,1}+a_{ij}^{V}})$ (7)
@@
While this technique is computationally expensive at large token counts, our models have a context length of 64, making the additional cost negligible.
III Experimental Setup
Our training data consists of self-play games generated with the AlphaZero process [1]. However, unlike that work, which worked in the reinforcement learning setting, we create a static dataset of self-play games from an older reinforcement learning run. That run used a transformer model of roughly 100 million parameters at 600 nodes per move which was almost identical in architecture to our main runs, barring its use of a slightly weaker position representation than that of our final architecture. We therefore work in the supervised setting without the need for online data generation, allowing for much faster training. Initial experiments showed no quality degradation on a static dataset.
All models we train employ a standard encoder-only backbone with a context of 64 tokens corresponding to the squares on the chessboard, with one of the position representations described in Section II. The training targets consist of the main policy and value targets in addition to a mix of auxiliary targets, and we use the Nadam optimzer with $\beta_{1}=0.9$, $\beta=0.98$, $\epsilon=10^{-7}$, and gradient clipping 10. The final model checkpoints used for evaluation are generated with stochastic weight averaging [6]. A more detailed description of the architecture and losses is provided in the Appendix.
@@ III Experimental Setup
Our largest model, CF-240M, has 15 encoder layers with an embedding depth of 1024, a head count of 32, and a feedforward depth of 4096, for a total of 243 million parameters. It was trained for 3.7 million steps with a batch size of 4096 on 8 A100 GPUs using data parallelism. The learning rate was initialized at $1*10^{-3}$ and manually reduced to $3*10^{-4}$ at 3.2 million steps and $1*10^{-4}$ at 3.6 million steps. The dataset consisted of 500 million games generated from mid-2023 to mid-2024, with each game containing roughly 200 positions.
@@
We also train a smaller model, CF-6M, with 8 encoder layers, an embedding depth of 256, a head count of 8, and a feedforward depth of 256, for a total of 6 million parameters. The same configuration is used for ablations of the position representation. CF-6M and its ablations were each trained on a single A100 GPU with a batch size of 2048. The learning rate was initialized at $5*10^{-4}$ and reduced to $1.58*10^{-4}$ and $5*10^{-5}$ at 1.6 and 1.8 million steps, respectively. The dataset consisted of 53 million games generated in April 2024. None of our training runs exhibited overfitting, which is likely due to the size of the datasets relative to the parameter counts of the models.
IV Results
@@
We compare the playing strength and puzzle-solving ability of agents constructed from our models to prior work and ablate the position representation in the attention mechanism to demonstrate the criticality of accurately modeling positional information.
@@
We include in our analysis two types of agents: those that function based on policy information by picking the move which is ranked highest in the policy vector predicted by the model, and those that function based on value information, emulating a search of depth 1 by evaluating the model for each legal move and selecting the move that maximizes the position evaluation. The policy strategy requires a single model evaluation, while the value maximization strategy requires an evaluation for each legal move. To estimate the floating point operations per evaluation (FLOPS) used by an agent of the value maximization type, we multiply the model FLOPS by 20, which is a rough estimate of the average number of legal moves available in a position.
@@
We construct agents from our models with both strategies, denoting an agent by its model name followed by the strategy it uses (e.g., CF-240M-policy and CF-240M-value). Policy and value agents are constructed from the AlphaZero model in the same way. Our analysis also includes models from Ruoss et al. [7] which use the value maximization approach. We use the final checkpoints of their main runs having parameter counts of 9 million, 136 million, and 270 million, referring to these as GC-9M, GC-136M, and GC-270M, respectively.
IV-A Playing Strength
@@
To estimate the playing strength of our agents, we play 1,000 games with them against the GC-270M agent of Ruoss et al. [7], tying the Elo rating of GC-270M to the value reported in that paper. We use the same opening book and testing configuration as that work to achieve a similar comparison to theirs. Elo values for the GC and AlphaZero agents are taken from that paper. The results are shown in LABEL:table:main and Fig. 1.
IV-B Puzzles
@@
We also evaluate the puzzle-solving ability of our agents on the puzzle database curated by Ruoss et al. [7]. Our evaluation follows the same methodology as that paper, with an agent deemed to solve a puzzle only if it correctly picks each move in the sequence of the solution. The results are shown in LABEL:table:main. Fig. 2 graphs the accuracy of our two CF-240M agents on this test set against the GC-136M and GC-270M agents of Ruoss et al. [7].
@@
Figure 2: Puzzle-solving ability on puzzles rated 1000-3000 of our CF-240M-policy and CF-240M-value agents against the GC-136M and GC-270M agents of Ruoss et al. [7].
IV-C Ablations
@@
We ablate the position representation of the CF-6M model, comparing the final test metrics of the three position representations described in Section II on a held-out test set of around 1 million games sharded from the training dataset. The reported statistics are for the main policy and value targets. As shown in LABEL:table:ablate, our chosen technique, that of Shaw et al. [4], substantially outperforms both the relative bias and absolute position embedding representations. At this parameter count, a doubling in model size yields an increase in policy accuracy of around 1.5%. The Shaw encoding outperforms the learned absolute position embedding by 1.83% policy accuracy, suggesting that a good choice of position representation can in large part replace the need for model scale.
IV-D Analysis
@@
The main results can be found in LABEL:table:main. Our agents consistently outperform prior work in both puzzle-solving ability and Elo performance at a fraction of the computational cost. Our CF-6M-policy agent outperforms the AlphaZero-policy agent in both metrics at 8x fewer FLOPS. At 30x fewer FLOPS, our CF-240M-policy agent matches the puzzle performance and exceeds the Elo performance of the grandmaster-level agent GC-270M of Ruoss et al. [7].
V Attention Maps
One advantage of transformers is the interpretability of their attention maps. We generate visualizations [^2] Code for generating attention visualizations is available at https://github.com/Ergodice/lc0-attention-visualizer. of the attention maps of our CF-240M model, which we plot in Fig. 3 and Fig. 4. We select a querying square and color squares based on their attention weights for the querying square. Small weights are colored purple and large weights are colored yellow.
The majority of attention heads in our models represent the movement of a particular piece, attending to squares which are a bishop’s, knight’s, rook’s, or king’s move away, depending on the specialization of the head. We give four examples from CF-240M in Fig. 3. From left to right then top to bottom, the heads shown are layer 5 head 24, layer 1 head 28, layer 1 head 14, and layer 15 head 22. The attention maps are fairly static across positions, always retaining their specialization even if the piece whose movement type they specialize in is no longer on the board.
Interspersed with these piece movement heads are several other heads with recognizable patterns. We give four examples in Fig. 4, describing them from left to right then top to bottom. Head 3 in layer 1 always attends to the opponent’s queen. Head 29 in layer 2 attends to squares which are of the same color on the checkerboard as the querying square. Head 14 in layer 15 attends to squares occupied by the opponent’s pieces with roughly equal weights. Finally, head 2 in layer 11 attends to pieces that can move to the querying square. This final type of head appears across model scales, though always in the second half of the encoder layers, suggesting that they rely on information about piece movement accumulated in early layers.
The variety of these miscellaneous heads suggests that the model develops a broad view of chess principles. The same-checkerboard-color head, for example, may provide information about the maneuverability of bishops. The head attending to the opponent’s pieces may be estimating the material difference between the players or gleaning the game stage, i.e., opening, middlegame, or endgame, from the types of pieces on the board.
Unlike in vision and language applications, proximity of tokens seems to have little effect on attention weights, confirming the importance of a position representation which can adapt to the topology of the chessboard. Another notable feature of the attention maps is that tokens rarely attend to themselves. This is in contrast to language models, where it is quite common [8].
Figure 3: Attention maps of heads corresponding to the movement of a particular piece. The square highlighted in red is the one producing the query.
Figure 4: Attention maps of several additional heads with easily interpretable patterns. The square highlighted in red is the one producing the query.
VI Playstyle
Position 1: fortress detection. CF-240M finds the only drawing move, Kg7, and evaluates the position as drawn.
Position 2: trapped piece detection. If white has a pawn on b2, CF-240M opts to trap the white rook on b3 by pushing b4. However, if the pawn on b2 is not present, it abandons this idea.
Position 3: long-term planning. CF-240M finds a famous king walk that was played out by Nigel Short.
Figure 5: Positions in which our models exhibit a humanlike understanding of the game, detecting positional ideas that elude top minimax-based engines.
Our largest model, CF-240M, appears to have an understanding of chess which is more akin to humans than that of conventional engines. In particular, it is capable of detecting difficult positional motifs that top chess engines struggle with. One position demonstrating this humanlike game understanding arose between Lc0 and Stockfish in Game 189 of the 23rd Computer Chess Championship Rapid Tournament, [^3] See https://www.chess.com/computer-chess-championship#event=ccc23-rapid-finals&game=189. where the former used a transformer with 190 million parameters trained by the authors. Lc0 managed to construct a fortress from a lost position, leading the game into a draw. A position from this game is shown as Position 1 in Fig. 5. CF-240M assigns the only drawing move, Kg7, a policy of 92.8%, and evaluates the position as drawn (2.2%W, 76.3%D, 21.5%L). Stockfish assigned the same position a +7 evaluation after searching 11 billion nodes, and its evaluation stayed above +3 for over 100 moves. Stockfish evaluations not directly interpretable as win probabilities, but a +3 evaluation corresponds roughly to an over 99% probability of winning.
Another position demonstrating this humanlike positional understanding arose between Lc0 and Stockfish in Game 163 of the 21st Computer Chess Championship Rapid Tournament, [^4] See https://www.chess.com/computer-chess-championship#event=ccc21-rapid-semifinals&game=163. where Lc0 used a transformer of around 100 million parameters trained by the authors. Shown as Position 2 in Fig. 5 is a critical moment in the game when Lc0 played b4, trapping the white rook on b3. CF-240M assigns this move the highest policy at 37% and evaluates the position as slightly winning (40.3%W, 54.1%D, 5.6%L). Lc0 ended up winning the game, with Stockfish’s evaluation only going negative several moves later despite searching hundreds of millions of positions.
If the pawn on b2 is removed and this trapping idea is no longer feasible, the evaluation becomes much more drawish (17.1%W, 77.7%D, 5.2%L), despite black gaining a material advantage. Additionally, the model loses interest in b4, with the policy for that move decreasing to 4.76% and Rc7 becoming the favored move.
The positional understanding of our CF-240M model seems to extend to long-term planning as well. As shown in Position 3 in Fig. 5, CF-240M picks the right move at each of white’s turns in a famous king walk played by GM Nigel Short as white against GM Jan Timman. The line continues 31. Kh2 Rc8 32. Kg3 Rce8 33. Kf4 Bc8 34. Kg5.
VII Related Work